OpenStudy (anonymous):

Find the indicated outputs for F(x)=4x^2-4x F(0)= F(-1)= F(2)=

OpenStudy (anonymous):

f(0) = 0 - 0 = 0 f(-1) = 4(-1)^2 -4(-1) = 4+4 = 8 f(2) = 4(2)^2 - 4(2) = 12

OpenStudy (saifoo.khan):

Insert the values instead of x.